OpenStudy (anonymous):

Juliet calculated her age to be eleven and five twelfths years. What number is equivalent to eleven and five twelfths? eleven point four one six eleven point four one six with a bar over four one six eleven point four one six with a bar over the six eleven point four one six with a bar over one six @Gatorgirl

OpenStudy (anonymous):

help

OpenStudy (campbell_st):

here is an online calculator just go 5 divide 12 http://web2.0calc.com/

OpenStudy (gatorgirl):

11.4166 years old

OpenStudy (gatorgirl):

that depends on what you mean by "equivalent" it is possible to re-write 11 5/12 as (132 + 5)/12 = 137/12.

OpenStudy (gatorgirl):

@travisbrown372 @KattGirl14 Am I right?

OpenStudy (anonymous):

r u sure

OpenStudy (travisbrown372):

yes 11 years and 5 months

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

wth no bar

OpenStudy (travisbrown372):

no bar yes

OpenStudy (anonymous):

A bag of fruit weighs zero point eight four kilograms. How can you write zero point eight four as a fraction in simplest form? eight and two-fifths eighty-four hundredths twenty-one twenty-fifths forty-two fiftieths

OpenStudy (gatorgirl):

C. Twenty-one twenty-fifths

OpenStudy (anonymous):

r u sure the only reason y im asking is cuz u said ur not good at math no offence me either

OpenStudy (anonymous):

C

OpenStudy (campbell_st):

5/12 = 0.4166666666666666666666666666666666666666666....... so its  I hope that helps...

OpenStudy (anonymous):

Annette found the decimal equivalent of a number using her calculator. A calculator shows the number zero point zero zero zero four four four four. How should Annette write the answer? zero point zero four with a bar over the zero four zero point zero zero zero four with a bar over zero zero zero four zero point zero zero zero four with a bar over four zero point zero zero zero four four four with a bar over zero zero zero four four four

OpenStudy (travisbrown372):

5/12 = 0.4166666666666666666666666666667

OpenStudy (travisbrown372):

no bar

OpenStudy (anonymous):

that does not help cuz not one of the answers

OpenStudy (campbell_st):

@travisbrown372... the 6 is recurring, the reason the calculator shows a 7 at the end is because it run our of display space and rounded the last 6 to a 7.

OpenStudy (jdoe0001):

\(\bf 11\cfrac{5}{12}\implies \cfrac{11\times 12+5}{12}\)

OpenStudy (gatorgirl):

zero point zero zero zero four with a bar over four

OpenStudy (gatorgirl):

next question

OpenStudy (campbell_st):

well the bar can only be used is the number repeats forever, called a recurring decimal.... and in the question the 4 only repeats for 4 times.... so its called a terminating decimal... so you need to write the full number 0.0004444

OpenStudy (anonymous):

Which is a requirement for a number to a rational number? The number can be written as a fraction. The denominator is an integer. The numerator is an integer. All of the above.

OpenStudy (gatorgirl):

D. All of the above.

OpenStudy (anonymous):

What decimal is equivalent to seven and 3 over two hundred fifty? seven point one two seven point zero one two seven point three two five seven point three zero two five

OpenStudy (gatorgirl):

7.012

OpenStudy (gatorgirl):

@i_need_help123 Medal please??

OpenStudy (anonymous):

Which number is equivalent to negative seven and one-third? negative seven point one three negative seven point one three with a bar over the three negative seven point three negative seven point three with a bar over the three

OpenStudy (travisbrown372):

https://www.mathway.com/ use this for algebra

OpenStudy (anonymous):

doesent work I tried

OpenStudy (gatorgirl):

-7.3 with bar over 3

OpenStudy (gatorgirl):

@travisbrown372 Medal please?